OpenStudy (anonymous):

what is the equation for a line that passes through the points (-4,-4) and (8, 17) ?

OpenStudy (anonymous):

-21/-12 7/4 -4=-4(7)/4 +b -4=-7+b 3=b y=(7/4)x+3

OpenStudy (anonymous):

thanks! do you know.... What is the slope of the line given by the following equation? y = -6x + 4